# PARAPHRASUS: A Comprehensive Benchmark for Evaluating Paraphrase Detection Models

### v1

@@ Abstract
The task of determining whether two texts are paraphrases has long been a challenge in NLP. However, the prevailing notion of paraphrase is often quite simplistic, offering only a limited view of the vast spectrum of paraphrase phenomena. Indeed, we find that evaluating models in a paraphrase dataset can leave uncertainty about their true semantic understanding. To alleviate this, we release paraphrasus  [^1] https://github.com/Andrian0s/paraphrasus a benchmark designed for multi-dimensional assessment of paraphrase detection models and finer model selection. We find that paraphrase detection models under a fine-grained evaluation lens exhibit trade-offs that cannot be captured through a single classification dataset.
@@ 1 Introduction
Our study was set in motion by a serendipitous finding. Like many other researchers at the time, we were benchmarking large language models (LLMs). We had particular interest in the paraphrasing detection task and evaluated LLama and others LLMs (Dubey et al., 2024; Alves et al., 2024) on the paraphrase challenge dataset PAWS-X (Yang et al., 2019). We consider paraphrasing as an important and challenging task, hand in hand with Quintilianus (95, p. 117) who lived 2,000 years ago, and of course, with many papers presented at CL and ML conferences such as Bhagat and Hovy (2013); Zhou and Bhat (2021); Krishna et al. (2023). Interestingly, LLMs appear to under-perform much on PAWS-X. Their performance lies slightly above random coin flip, and they are vastly outperformed by smaller BERT-based models fine-tuned on the training split of PAWS-X.
Figure 1: Percentage of paraphrases predicted the Semantic Text Similarity Dataset (STSBenchmark) dataset (Cer et al., 2017), binned by scores from 0 (completely dissimilar) to 5 (completely equivalent).
@@
    Within our benchmark, we also contribute two novel datasets. i) We annotate a test set of 338 semantically similar sentence pairs for paraphrase classification, creating a challenging non-adversarial test set that paraphrase detection methods plateau at 60% accuracy. ii) We exploit guidelines for Abstract Meaning Representation annotations, extracting expert examples of true paraphrases.
@@ 2 Related Work
Clearly, there are many NLP applications where paraphrasing and understanding paraphrases plays a crucial role (Mallinson et al., 2017). Some of these are semantic search (Reimers and Gurevych, 2019), style transfer (Krishna et al., 2020), machine translation (Madnani et al., 2007), and plagiarism detection (Barrón-Cedeño et al., 2013; Sharjeel et al., 2016). The evaluation of text generation is often interested in whether a candidate and the reference are paraphrases (e.g., Freitag et al., 2020; Thompson and Post, 2020; Nawrath et al., 2024).
@@
Recently, the PAWS Wiki dataset (Zhang et al., 2019) was introduced, providing an adversarial dataset for paraphrase classification through the application of word scrambling. They create challenging pairs by controlled word swapping and back-translation, followed by human quality check for fluency and paraphrase semantic. The dataset is split in 49,401 training sentence pairs and 8,000 test sentence pairs. pawsx (Yang et al., 2019) takes 2000 of the PAWS Wiki test samples and creates a 7-way multilingual test set in German, Spanish, French, Chinese, Japanese, and Korean using human translators. While these minimal pairs are interesting to perform a focused evaluation, its adversarial mode of dataset creation can introduce limitations on the coverage of linguistic formulation variation and hinder generalization abilities.
@@ 3 Proposed Benchmark
Paraphrases come in different flavors, and distinguishing types of paraphrases can be fuzzy (Bhagat and Hovy, 2013). Our benchmark therefore adopts an empirical approach to evaluating paraphrase detection models, according to three desiderata: i) It should reflect a broad spectrum of domains, to lower the impact of any superficial training that may have been acquired by a model. ii) It should reflect a broad spectrum of different flavors of paraphrases, e.g., varying strictness levels of paraphrases, negative examples with contradiction or neutral semantic relation. iii) It should help us select models for different paraphrase tasks, or select a single model that performs consistently across multiple categories. Table 1 gives an overview of the size and label distribution of each of the ten evaluation sets. More statistics are available in Appendix B.
@@ 3.1 The data: Ten Parts with Three Objectives
Our benchmark consists of 10 parts, with eight of them repurposed from various types of NLP tasks, such as natural language inference (NLI) and meaning representation annotation guidelines. By distributing the data across different domains, we aim to capture a wide range of paraphrase phenomena. An additional benefit of repurposing existing data could be that it helps alleviate concerns that the LLMs being evaluated may have incorporated this data during training. In addition to the repurposed data, two datasets are genuine paraphrase datasets, one of which is a human-annotated set of particularly challenging examples created by us.
The benchmark is divided into three objectives, along which we structure its description:
Classify paraphrases!
@@
  • mrpc: This is one of the first paraphrase datasets that has been released (Dolan and Brockett, 2005).
@@
  • sts-h: We propose using challenging sentence pairs without relying on adversarial strategies. To achieve this, we select highly similar sentence pairs from the STS dataset, which are annotated on a 5-point Likert scale. Pairs that are highly similar or paraphrased are situated at the upper end of this scale. We select all pairs from 4-5 and carefully create a high-quality set of 338 paraphrase annotations.
@@
    A semantics expert and a student annotated this subset independently, resulting in moderate-to-good Kappa of 0.63. Finally, disagreements were discussed and adjudicated. For the few cases where disagreements could not be resolved (13 out of 56 total), we assigned the non-paraphrase class. This decision was based on the reasoning that it is generally easier to argue that two sentences are not paraphrases than to prove the opposite.
Minimize paraphrase detection!
@@
In a dataset that contains no paraphrases, the primary objective of the model is to minimize the prediction of paraphrases. However, to ensure the task remains challenging, the paired texts should exhibit some degree of similarity. In this study, we introduce the following five datasets.
@@
  • snli, anli, xnli— Natural Language Inference (NLI) repurposed: NLI asks whether a hypothesis follows a given premise, labeling the relationship with entailed, neutral, or contradiction. We select those pairs that either stand in a neutral relationship, or in a contradiction. Since typically the premise is longer than the hypothesis, we control this bias by adding the flipped pairs. The resulting data are denoted by snli, sampled from the first large-scale NLI dataset (Bowman et al., 2015); anli taken from an adversarial version of the task (Nie et al., 2020); and the cross-lingual xnli Conneau et al. (2018) that allows us to study paraphrase detection models in the minimization task in different languages.
@@
  • sts, sick— Similarity repurposed: Orthogonally to how we created the sts-h data by using only extremely similar pairs, we now select only pairs that we know that they are not paraphrases. To ensure maximum data quality and lower the possibility of annotation confusions in the upper spectrum of the similarity Likert scale, we select only values 1-3 as negative pairs, excluding the range of 3 to 5. The resulting data are gathered from two datasets (Marelli et al., 2014; Cer et al., 2017).
Maximize paraphrase detection!
@@
If we know that a dataset contains only pairs of paraphrases, the natural expectation towards a model would be to maximize its paraphrase detection rate. We created two datasets to test this behavior:
@@
  • true: We create a dataset of simple, guaranteed paraphrases. The objective for a paraphrase model using this dataset is to correctly identify as many paraphrases as possible. To ensure the simplicity and accuracy of these examples, we leverage Abstract Meaning Representation annotation guidelines (Banarescu et al., 2013). [^2] https://github.com/amrisi/amr-guidelines AMR has the goal of mapping the same meaning to the same structure, and highlights this in its annotation guidelines by presenting graphs together with a group of semantically equivalent (but structurally different) sentences. For every unique meaning graph, we extract the $n$-sized set of represented sentences and create $\frac{n^{2}-n}{2}$ of unique paraphrase pairs. [^3] Since AMR currently captures many, but not all semantic aspects (Sadeddine et al., 2024), we exclude any parts of the annotation guidelines where AMR conflates two aspectually divergent sentences into a single representation. As the guidelines succinctly treat diverse phenomena, there can be short phrases, like “20 km” and “20 kilometers”, but most of them are short sentences, e.g., “The boy desires the girl to believe him.” and “The boy has a desire to be believed by the girl.”  [^4] The license of the AMR guidelines allows us to publicly release this subset of true paraphrases.
@@
  • simp: From the AMR guidelines, we retrieved simple and accurate paraphrases; however, the manually constructed nature of the data limits its diversity. To create a more diverse dataset of paraphrases, albeit at the cost of introducing more noise, we repurpose text simplification sentence pairs from Vajjala and Lučić (2018)’s OneStopEnglish corpus.
3.2 Evaluation metrics
To achieve a final ranking of systems, we need to select a metric. While there are many choices for a metric, and each has their benefits and drawbacks (Opitz, 2024), we wish for an interpretable and robust measure. For each dataset, we calculate an error percentage. For the classification tasks, this is the ratio of wrong predictions. For minimization datasets, like the neutral and contradictory pairs of the NLI data, this is the ratio of positive predictions. For the maximization datasets, like the certified paraphrases in true, this is the ratio of negative predictions. To achieve a robust final score, we calculate an unweighted average over the three objectives of paraphrasus, and then another unweighted average over these averages.
4 Testing Paraphrase Detection Models
Sentence 1: {sentence1}
Sentence 2: {sentence2}
Answer with ’Yes’ or ’No’
Figure 2: For P1, P2, and P3, the paraphrase notions we ask for are “paraphrases”, “semantically equivalent” and “expressing the same content” respectively. For the ICL expanded prompt, see Appendix A
  1. 1.
  2. 2.
4.1 What can we learn from pawsx?
We replicate the approach from the original pawsx study by Yang et al. (2019) and fine-tune a multilingual encoder model, XLM-RoBERTa base (XLM-R) Conneau et al. (2020), on the English Paws Wiki training set. Using default fine-tuning hyperparameters, we train for 6 epochs. During initial evaluation, we observed considerable variance between different fine-tuning seeds on paraphrasus. To reduce the impact of this variance on our results, we report the average performance based on three high-performing checkpoints, each corresponding to a different seed, yielding a total of nine predictions.
4.2 What can we learn from LLMs?
Recent research has indicated that LLMs have developed a strong understanding of human language, which can be accessed through natural language prompts. In this study, we prompt LLama3 8B (Dubey et al., 2024) with three different paraphrase notions along with their in-context learning versions.
@@ 4.2 What can we learn from LLMs?
Results are error percentage Classify! Minimize! Maximize! Averages
@@
LLama3 Instruct P3 38.1 41.6 37.5 1.3 1.7 1.3 3.5 0.1 35.3 37.5 39.1 1.6 36.4 25.7
LLama3 Instruct P1-ICL_K4 39.0 44.7 33.2 2.0 2.0 1.4 3.5 0.3 29.9 33.3 38.9 1.8 31.6 24.1
LLama3 Instruct P2-ICL_K4 34.1 41.7 37.5 0.9 0.8 0.0 3.1 0.0 40.1 42.3 40.4 1.0 41.2 24.6
LLama3 Instruct P3-ICL_K4 33.2 39.1 47.1 0.5 0.8 0.1 2.4 0.0 50.9 45.5 39.8 0.7 48.2 24.4
Table 2: Main results on paraphrasus. For description of prompts P1, P2, P3, and their ICL variants, see §4.2.
@@
Given that multiple “personalities” can be elicited from LLMs (Chan et al., 2024), we leverage this property to design three distinct, minimalistic prompts that vary only in how the concept of “paraphrase” expressed. The first prompt (henceforth denoted as P1) is most straightforward, just asking the model to judge whether two sentences are paraphrases. The second prompt (P2) intends to emulate a person with background in semantics, asking the model whether the following two sentences are semantically equivalent. We speculate that this prompt triggers a model mode that very strictly judges the pairs, refraining from assigning a positive label if there is only a minor semantic difference. The third prompt (P3) aims to kindle a person who is primed from Information Retrieval: The model is asked whether the following two sentences are about the same content. We speculate that this prompt provides a suitable balance between the straightforward P1 prompt and the presumed strict prompting P2.
@@
Our second version of these prompt personas leverages the “emergent capability” of in-context learning (Brown et al., 2020), and is henceforth denoted by the (ICL_K4) suffix. Compared to the original, during inference time, we sample two positive and two negative sentence pairs from the training set of PAWS-Wiki and prepend them to the prompt.
5 Main Benchmark Results
The main results on paraphrasus are shown in Table 2.
What model is best overall?
Interestingly, our most simple LLM setup (P1) demonstrates the best overall result (21% errors on average) on paraphrasus, outperforming the second best model that uses in-context examples from pawsx (P1-ICL_K4) by 3 percentage points (pp). While the smaller fine-tuned XLM-R model excels on pawsx with an error rate of only 15.2% it struggles to generalize across the other objectives and ranks last overall, showing an overall error of 27.5%.
What model is best for specific objectives?
Findings on paraphrasus show that each model exhibits unique strengths. When examining the averages for each objective (right side of Table 2), LLama with prompt P3-ICL_K4 (Min!, 0.7% average error) outperforms others in the minimization objective, while LLama with prompt P1 (Max!, 11.8% average error) excels in the maximization objective. In the Classification objective, the trained XLM-R model achieves the best performance. However, it is important to note that this particular average is skewed due to the model’s training on pawsx. In the other two classification datasets, XLM-R is outperformed by LLama models, with up to a 15 percentage point error difference on sts-h (using prompt P2) and a 9.8 percentage point difference on mrpc (using prompt P1).
Zero Shot or In Context Learning
In the zero shot experiments, the different paraphrase notions in the prompt lead to different levels of strictness and overall result, whereas in the ICL experiments the three prompts behave and perform similarly. Adding in-context examples to (P1) seems to calibrate the model to be stricter and reduce false positives, but at the cost of a 19.8 percentage point error difference in the maximization average, and a lower overall result.
6 Discussion
Results are error percentage Classify! Minimize! Maximize! Averages
Model $\leftarrow$ PAWS-X % Negatives to EasyNegs.
@@ 6 Discussion
Table 3: Ablation study A: XLM-R fine-tuned on PAWS Wiki training samples where a percentage of the second sentences of the negative pairs was exchanged with a random sentence.
Results are error percentage Classify! Minimize! Maximize! Averages
Model $\leftarrow$ PAWS-X EN Train Labeled by Prompt
@@
[…]
Table 4: Ablation study B: XLM-R fine-tuned on PAWS Wiki training samples but instead using predicted LLama3 labels as targets (for description of prompts P1, P2, P3 see 4.2)

### v2

@@ Abstract
The task of determining whether two texts are paraphrases has long been a challenge in NLP. However, the prevailing notion of paraphrase is often quite simplistic, offering only a limited view of the vast spectrum of paraphrase phenomena. Indeed, we find that evaluating models in a paraphrase dataset can leave uncertainty about their true semantic understanding. To alleviate this, we create paraphrasus, a benchmark designed for multi-dimensional assessment, benchmarking and selection of paraphrase detection models. We find that paraphrase detection models under our fine-grained evaluation lens exhibit trade-offs that cannot be captured through a single classification dataset. Furthermore, paraphrasus allows prompt calibration for different use cases, tailoring LLM models to specific strictness levels. paraphrasus includes 3 challenges spanning over 10 datasets, including 8 repurposed and 2 newly annotated; we release it along with a benchmarking library at https://github.com/impresso/paraphrasus
@@ 1 Introduction
Our study was set in motion by a serendipitous finding. Like many other researchers at the time, we were benchmarking large language models (LLMs). We had particular interest in the paraphrasing detection task [^1] We consider paraphrasing as an important and challenging task, hand in hand with Quintilianus (95, p. 117) who lived 2,000 years ago, and of course, with many papers presented at CL and ML conferences such as Bhagat and Hovy (2013); Zhou and Bhat (2021); Krishna et al. (2023). and evaluated Llama and others LLMs (Dubey et al., 2024; Alves et al., 2024) on the adversarial paraphrase detection test set PAWS-X (Yang et al., 2019). Interestingly, LLMs appear to under-perform much on PAWS-X. Their performance lies slightly above random coin flip, and they are vastly outperformed by smaller BERT-based models fine-tuned on the training split of PAWS-EN.
Figure 1: Percentage of paraphrases predicted on the Semantic Text Similarity Dataset (STSBenchmark) dataset (Cer et al., 2017), binned by scores from 0 (completely dissimilar) to 5 (completely equivalent). Human annotation comes from the sts-h human annotation we perform.
@@
    Within our benchmark, we also contribute two novel datasets. i) We annotate a test set of 338 semantically similar sentence pairs for paraphrase classification, creating a challenging test set that paraphrase detection methods plateau at 60% accuracy. ii) We exploit guidelines for Abstract Meaning Representation annotations, extracting expert examples of true paraphrases.
@@ 2 Related Work
Clearly, there are many NLP applications where paraphrasing and understanding paraphrases plays a crucial role (Mallinson et al., 2017). Some of these are semantic search (Reimers and Gurevych, 2019), style transfer (Krishna et al., 2020), machine translation (Madnani et al., 2007), plagiarism (Barrón-Cedeño et al., 2013; Sharjeel et al., 2016), and text reuse (Büchler et al., 2012). The evaluation of text generation is often interested in whether a candidate and the reference are paraphrases (e.g., Freitag et al., 2020; Thompson and Post, 2020; Nawrath et al., 2024).
@@
Recently, the PAWS dataset (Zhang et al., 2019) was introduced, providing an adversarial dataset for paraphrase classification through the application of word scrambling. They create challenging pairs by controlled word swapping and back-translation, followed by human quality check for fluency and paraphrase semantic. The dataset is split in 49,401 training sentence pairs and 8,000 test sentence pairs. pawsx (Yang et al., 2019) takes 2000 of the PAWS Wiki test samples and creates a 7-way multilingual test set in German, Spanish, French, Chinese, Japanese, and Korean using human translators. While these minimal pairs are interesting to perform a focused evaluation, its adversarial mode of dataset creation can introduce limitations on the coverage of linguistic formulation variation and hinder generalization abilities.
@@ 3 Proposed Benchmark
Paraphrases come in different flavors, and distinguishing types of paraphrases can be fuzzy (Bhagat and Hovy, 2013). Our benchmark therefore adopts an empirical approach to evaluating paraphrase detection models, according to three desiderata: i) It should reflect a broad spectrum of domains, to lower the impact of any superficial training that may have been acquired by a model. ii) It should reflect a broad spectrum of different flavors of paraphrases, e.g., varying strictness levels of paraphrases, negative examples with contradiction or neutral semantic relation. iii) It should help us select models for different paraphrase challenges, or select a single model that performs consistently across multiple categories. Table 1 gives an overview of the size and label distribution of each of the ten evaluation sets. More statistics are available in Appendix C.
@@ 3.1 The data: Ten Parts with Three Objectives
Our benchmark consists of 10 parts, with eight of them repurposed from various types of NLP tasks, such as natural language inference (NLI) and meaning representation annotation guidelines. By distributing the data across different domains, we aim to capture a wide range of paraphrase phenomena. An additional benefit of repurposing existing data could be that it helps alleviate concerns that the LLMs being evaluated may have incorporated this data during training. In addition to the repurposed data, three datasets are genuine paraphrase datasets, one of which is a human-annotated set of particularly challenging examples created by us.
The benchmark is divided into three challenges, along which we structure its description:
Classify!
@@
  • mrpc: As also mentioned in Section 2, mrpc is one of the first paraphrase detection dataset that has been released (Dolan and Brockett, 2005).
@@
  • sts-h: We propose using challenging sentence pairs without relying on adversarial strategies. To achieve this, we select highly similar sentence pairs from the STS dataset, which are annotated on a 5-point Likert scale. Potential paraphrase candidates are situated at the upper end of this scale. We select all pairs from 4-5 and carefully create a high-quality set of 338 paraphrase annotations.
@@
    A semantics expert and a student annotated this subset independently, resulting in moderate-to-good Kappa of 0.63 Cohen (1960). Finally, disagreements were discussed and adjudicated. For the few cases where disagreements could not be resolved (13 out of 56 total), we assigned the non-paraphrase class. This decision was based on the reasoning that it is easier to argue that two sentences are not paraphrases than to prove the opposite.
Minimize positive predictions!
@@
In a dataset that contains no paraphrases, the primary objective of the model is to minimize the prediction of paraphrases. However, to ensure the task remains challenging, the paired texts should exhibit some degree of similarity. In this study, we repurpose the following five datasets.
@@
  • snli, anli, xnli— Natural Language Inference (NLI) repurposed: NLI asks whether a hypothesis follows a given premise, labeling the relationship with entailed, neutral, or contradiction. We select those pairs that either stand in a neutral relationship, or in a contradiction. Since typically the premise is longer than the hypothesis, we control this bias by adding the flipped pairs. The resulting data are denoted by snli, sampled from the first large-scale NLI dataset (Bowman et al., 2015); anli taken from an adversarial version of the task (Nie et al., 2020); and the cross-lingual xnli Conneau et al. (2018)  [^2] Limited to EN, DE, FR, ES, ZH due to compute costs. that allows us to study paraphrase detection models in the minimization task in different languages.
@@
  • sts, sick— Similarity repurposed: Orthogonally to how we created the sts-h data by using only extremely similar pairs, we now select only pairs that we know that they are not paraphrases. To ensure maximum data quality and lower the possibility of annotation confusions in the upper spectrum of the similarity Likert scale, we select only values 0-3 as negative pairs, excluding the range of 3 to 5. The resulting data are gathered from two datasets (Marelli et al., 2014; Cer et al., 2017).
Maximize positive predictions!
@@
If we know that a dataset contains only pairs of paraphrases, the natural expectation towards a model would be to maximize its paraphrase detection rate. We selected two datasets to test this behavior:
@@
  • true: We create a dataset of simple, guaranteed paraphrases. The objective for a paraphrase model using this dataset is to correctly identify as many paraphrases as possible. To ensure the accuracy of these examples, we leverage Abstract Meaning Representation annotation guidelines (Banarescu et al., 2013). [^3] https://github.com/amrisi/amr-guidelines AMR has the goal of mapping the same meaning to the same structure, and highlights this in its annotation guidelines by presenting graphs together with a group of semantically equivalent (but structurally different) sentences. For every meaning graph, we extract the $n$-sized set of example sentences and create $\frac{n^{2}-n}{2}$ true paraphrase pairs. [^4] Since AMR currently captures many, but not all semantic aspects (Sadeddine et al., 2024), we exclude any parts of the annotation guidelines where AMR conflates two aspectually divergent sentences into a single representation. As the guidelines succinctly treat diverse phenomena, there can be short phrases, like “20 km” and “20 kilometers”, but most of them are short sentences, e.g., “The boy desires the girl to believe him.” and “The boy has a desire to be believed by the girl.”  [^5] The license of the AMR guidelines allows us to publicly release this subset of true paraphrases.
@@
  • simp: From the AMR guidelines, we retrieved simple and accurate paraphrases; however, the manually constructed nature of the data limits its diversity. To create a more diverse dataset of paraphrases, albeit at the cost of introducing more noise, we repurpose parallel sentence pairs of different readability levels (Advanced, Intermediate, Elementary) from Vajjala and Lučić (2018)’s OneStopEnglish.
@@
Results are error percentages $\boldsymbol{\downarrow}$ Classify! Minimize! Maximize! Averages
@@
Llama3 Instruct P3 38.1 41.7 37.5 1.3 1.7 1.3 3.5 0.0 35.3 37.5 39.1 1.6 36.4 25.7
Llama3 Instruct P1-ICL_K4 39.0 44.7 33.2 1.9 2.0 2.8 3.5 0.3 29.9 33.3 39.0 2.1 31.6 24.2
Llama3 Instruct P2-ICL_K4 34.1 41.7 45.2 0.8 0.8 0.3 3.1 0.0 40.1 42.3 40.3 1.0 41.2 27.5
Llama3 Instruct P3-ICL_K4 33.2 39.1 46.7 0.5 0.8 0.3 2.4 0.0 50.9 45.5 39.7 0.8 48.2 29.6
Table 2: Main results on paraphrasus. For description of prompts P1, P2, P3, and their ICL variants, see §4.2. All results are error percentages: Lower numbers are better. The classification method with the best overall performance is marked with an asterisk ().
3.2 Evaluation Metric
To achieve a final ranking of systems, we need to select a metric. While there are many choices for a metric, and each has their benefits and drawbacks (Opitz, 2024), we wish for an interpretable and robust measure. For each dataset, we calculate a classification $error$ percentage. For datasets from the classify challenge, this is the ratio of wrong predictions. For datasets where the goal is to minimize the amount of positive predictions, like the neutral and contradictory pairs of the NLI data, this is the ratio of positive predictions (in that case, equivalent to false positive rate). For the datasets in the maximization challenge, such as true, this is the ratio of negative predictions (equivalent to false negative rate). To achieve a robust final score, we calculate an unweighted average per each of the three challenges, and then another unweighted average over these averages:
$\overline{Err}=\frac{\sum\limits_{\text{challenge}\in\text{benchmark}}\left(\frac{\sum\limits_{\text{dataset}\in\text{challenge}}\text{error(dataset)}}{|\text{challenge}|}\right)}{|\text{benchmark}|}$
The $\overline{Err}$ models the probability of observing a wrong prediction, given we randomly select one of the three challenges, and randomly pick an example from a random dataset of the task. Since all scores on paraphrasus are error measurements (or their averages), lower numbers are better.
4 Testing Paraphrase Detection Models
Sentence 1: {sentence1}
Sentence 2: {sentence2}
Answer with ’Yes’ or ’No’
Figure 2: For P1, P2, and P3, the paraphrase notions we ask for are “paraphrases”, “semantically equivalent” and “expressing the same content” respectively. For the ICL expanded prompt, see Appendix A.
  1. 1.
  2. 2.
4.1 What can we learn from pawsx?
We replicate the approach from the original pawsx study by Yang et al. (2019) and fine-tune a multilingual encoder, XLM-RoBERTa base (XLM-R) Conneau et al. (2020) on the PAWS Wiki English training set. Using default fine-tuning hyperparameters, we train for 6 epochs. During initial evaluation, we observed considerable variance between different fine-tuning seeds on paraphrasus. To reduce the impact of this variance on our results, we report the average performance based on three high-performing checkpoints, each corresponding to a different seed, yielding a total of nine predictions.
4.2 What can we learn from LLMs?
Recent research has indicated that LLMs have developed a strong understanding of human language, which can be accessed through natural language prompts. In this study, we prompt Llama3 Instruct 8B (Dubey et al., 2024) quantized to 4 bits with three different paraphrase notions along with their in-context learning equivalents.
@@ 4.2 What can we learn from LLMs?
Given that multiple “personalities” can be elicited from LLMs (Chan et al., 2024), we leverage this property to design three distinct, minimalistic prompts that vary only in how the concept of “paraphrase” expressed. The first prompt (henceforth denoted as P1) is most straightforward, just asking the model to judge whether two sentences are paraphrases. The second prompt P2 intends to emulate a person with background in semantics, asking the model whether the following two sentences are semantically equivalent. We speculate that this prompt triggers a model mode that very strictly judges the pairs, refraining from assigning a positive label if there is only a minor semantic difference. The third prompt P3 aims to kindle a person who is primed from Information Retrieval: The model is asked whether the following two sentences are about the same content. We speculate that this prompt provides a suitable balance between the straightforward P1 prompt and the presumed strict prompting P2.
@@
Our second version of these prompt personas leverages the “emergent capability” of in-context learning (Brown et al., 2020), and is henceforth denoted by the ICL_K4 suffix. Compared to the original, during inference time, we sample two positive and two negative sentence pairs from the training set of PAWS-Wiki and prepend them to the prompt.
Results are error percentages $\boldsymbol{\downarrow}$ Classify! Minimize! Maximize! Averages
Model $\leftarrow$ PAWS-EN % Negatives to EasyNegs.
@@
Table 3: Ablation study A: XLM-R fine-tuned on PAWS EN Wiki training samples where a percentage of the second sentences of the negative pairs was exchanged with a random sentence. All results are error percentages: Lower numbers are better. The classification method with the best overall performance is marked with an asterisk ().
Results are error percentages $\boldsymbol{\downarrow}$ Classify! Minimize! Maximize! Averages
Model $\leftarrow$ PAWS EN Train Labeled by Prompt
@@
XLM-R $\leftarrow$ Llama3 $\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}P1}\land{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}P2}\land{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}P3}}$ 28.4 54.6 34.0 33.7 49.0 55.7 40.6 19.0 38.1 10.7 39.0 39.6 24.4 34.3
Table 4: Ablation study B: XLM-R fine-tuned on PAWS Wiki training samples but instead using predicted Llama3 labels as targets (for description of prompts P1, P2, P3 see 4.2). $X$land $Y$: The positive label is assigned if and only if it is predicted by both prompt $X$ and $Y$. All results are error percentages: Lower numbers are better. The classification method with the best overall performance is marked with an asterisk ()
5 Main Benchmark Results
The main results on paraphrasus are shown in Table 2.
What model is best overall?
Interestingly, our most simple LLM setup (P1) demonstrates the best overall result (21% errors on average) on paraphrasus, outperforming the second best model that uses in-context examples from pawsx (P1-ICL_K4) by 3.2 percentage points (pp). While the smaller fine-tuned XLM-R model excels on pawsx with an error rate of only 15.2% it struggles to generalize across the other objectives and ranks second to last, showing an overall error of 27.5%.
What model is best for specific objectives?
Findings on paraphrasus show that each model exhibits unique strengths. When examining the averages for each objective (right side of Table 2), Llama with prompt P3-ICL_K4 (Min!, 0.8% average error) outperforms others in the minimization objective, while Llama3 with prompt P1 (Max!, 11.8% average error) excels in the maximization objective. In the classification objective, the trained XLM-R model achieves the best performance. However, it is important to note that this particular average is skewed due to the model’s training on pawsx. In the other two classification datasets, XLM-R is outperformed by Llama3 prompts, with up to a 14.8 pp error difference on sts-h (using prompt P2) and a 9.8 pp difference on mrpc (using prompt P1).
Zero Shot or In Context Learning
In the zero-shot experiments, the different paraphrase notions in the prompt lead to different levels of strictness and overall result, whereas in the ICL experiments the three prompts behave stricter. Adding in-context examples to P1 seems to calibrate the model to reduce false positives, but at the cost of a 19.8 pp error difference in the maximization average, and a lower overall result.
6 Discussion